# Power allocation for cell-free MIMO integrated sensing and communication

### v1

@@ Abstract
In this paper, we investigate integrated sensing and communication (ISAC) in a cell-free (CF) multiple-input multiple-output (MIMO) network with single-antenna access points (APs), where each AP functions either as a transmitter for both sensing and communication or as a receiver for target-reflected signals. We derive closed-form Cramér–Rao lower bounds (CRLBs) for location and velocity estimation under arbitrary power allocation ratios, assuming the radar cross-section (RCS) is deterministic and unknown over the observation interval. A power allocation optimization problem is formulated to maximize the communication signal-to-interference-plus-noise ratio (SINR), subject to CRLB-based sensing constraints and per-transmitter power limits. To solve the resulting nonlinear and non-convex problem, we propose a penalty function and projection-based modified conjugate gradient algorithm with inexact line search (PP-MCG-ILS), and an alternative method based on a modified steepest descent approach (PP-MSD-ILS). Additionally, for power minimization in pure sensing scenarios, we introduce a penalty function-based normalized conjugate gradient algorithm (P-NCG-ILS). We analyze the convergence behavior and qualitatively compare the computational complexity of the proposed algorithms. Simulation results confirm the accuracy of the derived CRLBs and demonstrate the effectiveness of the proposed power allocation strategies in enhancing both sensing and overall ISAC performance.
@@ I Introduction
Integrated sensing and communication (ISAC) has emerged as a key enabling technology for sixth-generation (6G) applications, such as smart cities, autonomous driving, and remote healthcare [1, 2]. Unlike traditional systems with separate designs for communication and sensing, ISAC unifies these functionalities within shared hardware platforms and spectral resources, thereby significantly enhancing spectral, energy, and hardware efficiency [3, 4].
Depending on the geometric deployment of transmitters and receivers, various ISAC system architectures have been proposed. In mono-static ISAC systems—where the transmitter and receiver are co-located—advanced beamforming techniques are employed to mitigate self-interference in full-duplex operations [5, 6]. In contrast, cell-free (CF) or multi-static ISAC systems involve widely distributed transmitters and receivers. These systems inherit advantages from both CF communication and distributed multi-static radar due to their architectural similarities [7]. MIMO (multiple-input, multiple-output) technology plays a central role by employing orthogonal waveforms across transmitters to achieve multiplexing and diversity gains for both communication and sensing [8]. Owing to these benefits, CF MIMO-empowered ISAC has been actively explored for unified communication-sensing operations [9, 10, 11, 12, 13, 14].
@@
ISAC systems address a range of sensing tasks, including target detection [15, 16, 17], and parameter estimation for radar cross-section (RCS) [18, 19], location [20, 21, 22], velocity [23, 24], and angle [25, 26]. These capabilities support diverse applications from navigation to environmental monitoring. A fundamental challenge in ISAC system design is the formulation of robust performance metrics, which are critical for resource optimization and balancing communication-sensing trade-offs [3, 4]. While unified metrics based on mutual information [27] and Kullback–Leibler divergence [28] have been proposed, tailoring such metrics to specific sensing objectives remains challenging. Consequently, task-specific metrics for sensing and communication continue to be widely adopted.
For communication, Shannon capacity underpins metrics such as spectral efficiency [29], achievable rate [30], and signal-to-interference-plus-noise ratio (SINR). Other metrics like coverage probability and energy efficiency have also been addressed in scenarios with stringent requirements [31]. On the sensing side, metrics such as sensing mutual information [32, 33], sensing SINR [9, 10], and the Cramér–Rao lower bound (CRLB) [34, 35, 36, 37, 38, 39, 40, 41] have been explored extensively, with CRLB-based designs proving particularly effective.
CRLB-based approaches fall into two categories depending on the RCS model:
Random RCS assumption: RCS values are treated as random variables, and likelihood functions are derived by averaging over their distributions [34, 35, 36, 37]. These studies established CRLBs for joint location and velocity estimation, and later extended them to multi-target scenarios.
Deterministic RCS assumption: RCS values are assumed to be deterministic but unknown and spatially diverse [38, 40, 41, 39]. This model is particularly relevant in slowly varying environments or when prior RCS knowledge is unavailable. Under this setting, CRLBs for both single-target and multiple-target location estimation have been analyzed, with [39] showing convergence to single-target performance as target separations increase.
@@
However, most of the above CRLB formulations rely on continuous-time signal models, while practical radar systems process discrete-time signals, where the sampling rate significantly impacts sensing accuracy. Our prior work [42] introduces a discrete-time likelihood function over Gaussian channels as a more realistic model. Building on this, we derive CRLBs for joint location and velocity estimation under the deterministic RCS assumption in both single- and multiple-target scenarios.
@@
Another critical challenge lies in resource allocation, involving time [43], frequency [44, 45], space [46, 47], and power [9, 10, 12, 11, 13, 14]. While recent works have addressed power allocation for maximizing sensing SINR or minimizing power under CRLB and communication constraints [9, 10, 12, 11, 13], they typically consider localization only, neglecting velocity or angular estimation, and are based on continuous-time models that overlook the impact of sampling rate.
In contrast, this paper focuses on joint location and velocity estimation using discrete-time signal models. Even with simplified CRLBs, the resulting CRLB-constrained power allocation problems are highly nonlinear and non-convex, posing significant challenges to analytical optimization. To address this, we leverage advanced optimization techniques including penalty methods, projection operations, and gradient-based search, building algorithms suitable for such complex constrained problems.
I-A Contribution
This paper addresses the power allocation problem in a CF MIMO-ISAC system with widely distributed single-antenna access points (APs), where each AP operates as either an ISAC transmitter or sensing receiver. We derive CRLBs for joint location and velocity estimation using discrete received signals under the deterministic RCS assumption. These CRLBs are adopted as sensing metrics to guide power allocation under communication SINR maximization objectives. The main contributions are as follows.
1) We derive the approximate CRLBs for joint location and velocity estimation of a single target under arbitrary power allocations. The CRLBs are validated against MLE results, confirming their effectiveness as performance metrics. With the sensing requirements constrained by the CRLBs of both location and velocity estimation, we propose a Penalty-function based Normalized Modified Conjugate Gradient algorithm with Inexact Line Search (P-NCG-ILS) to generate a power allocation scheme for the total power minimization. Simulation results verify its effectiveness.
@@ I-A Contribution
2) We formulate the CF MIMO-ISAC power allocation problem as a communication SINR maximization task, subject to constraints on both location and velocity CRLBs and total/individual power budgets. To tackle the non-convexity and nonlinearity, we propose a Penalty-function and Projection-based Modified Conjugate Gradient algorithm with Inexact Line Search (PP-MCG-ILS) together with PP-MSD-ILS algorithm based on the modified steepest descent algorithm (MSD).
@@
3) The proposed PA algorithms for sensing and CF MIMO-ISAC can be broadly applicable to constrained non-convex optimization problems involving affine equality constraints (handled via projection) and nonlinear inequality constraints (handled via penalties). We analyze their convergence and computational complexity. Simulation results demonstrate the robust performance, fast convergence, and effectiveness of the proposed algorithms (e.g., PP-MCG-ILS) in optimizing joint sensing and communication performance.
@@
The remainder of the following parts of this paper is organized as follows. Section II describes the received signal model for ISAC. Section IV formulates the power allocation problem for ISAC. Section V details the proposed ISAC power allocation algorithms. Section VII gives the convergence and computational complexity analysis. Section VIII illustrates the simulation results. Section IX concludes this paper.
@@ II Signal Model
As shown in Fig. 1, we consider a CF MIMO-ISAC system with widely separated APs. Each AP either serves as an ISAC transmitter or a sensing receiver, similar to the setup in [9, 10]. Assume there are $N$ single-antenna transmitters and $K$ single-antenna receivers, with the positions of the transmitter $n$ and receiver $k$ denoted as ${\bm{l}}_{n}=[x_{n},y_{n}]^{\rm T}$ and ${\bm{l}}_{k}=[x_{k},y_{k}]^{\rm T}$, respectively. All transmitters and receivers are scheduled and managed by the edge cloud and the core network, e.g., the fifth generation (5G) communication networks. Denote the location and velocity vectors of the $q$th sensing target as ${\bm{l}}_{q}=[x_{q},y_{q}]^{\rm T}$ and ${\bm{v}}_{q}=[v_{x,q},v_{y,q}]^{\rm T}$, respectively.
@@
The Doppler frequency intorduced by the relative radial velocity of target $q$ versus any transmitter $n$ and receiver $k$, is computed as
$\small f_{n,k,q}=\dfrac{1}{\lambda}{\bm{v}}_{q}^{\rm T}\left(\frac{{\bm{l}}_{n}-{\bm{l}}_{q}}{\|{\bm{l}}_{n}-{\bm{l}}_{q}\|_{2}}+\frac{{\bm{l}}_{k}-{\bm{l}}_{q}}{\|{\bm{l}}_{k}-{\bm{l}}_{q}\|_{2}}\right),$ (1)
where $\lambda$ denotes the wavelength. Denote $\tau_{n,k,q}$ as the propagation delay between the $n$th transmitter and the $k$th receiver, as reflected by the $q$th target. Then, we have
$\small\tau_{n,k,q}$ $=\tau_{n,q}+\tau_{k,q},$ (2a)
$\tau_{n,q}$ $={\|{\bm{l}}_{n}-{\bm{l}}_{q}\|_{2}}/{c},$ (2b)
$\tau_{k,q}$ $={\|{\bm{l}}_{k}-{\bm{l}}_{q}\|_{2}}/{c},$ (2c)
where $\tau_{n,q}$ and $\tau_{q,k}$ represent the propagation delays from the $n$th transmitter to the $q$th target, and from the target to the $k$th receiver, respectively, and $c$ denotes the speed of light. In this paper, we consider the single-target scenario, and thus the subscript $q$ can be omitted.
@@
where $P$ denotes the total available energy, $\rho_{n}$ is the power allocation (PA) factor with $\sum_{n=1}^{N}\rho_{n}=1$, ${b}_{n}$ is the normalized transmit beamforming weight with $|{b}_{n}|^{2}=1$, and $\varsigma$ is the complex data symbol with $\mathbb{E}\{|\varsigma|^{2}\}=1$. Denote ${\bm{\rho}}=[\rho_{1},\rho_{2},\cdots,\rho_{N}]^{\rm T}$ by the vector of the PA coefficient of the $N$ transmitters.
For target-reflected reception, we refer the term “RCS” as encompassing both channel fading and radar RCS. Then, we define the RCS as ${\alpha}_{n,k}$ over transmitter $n$ and receiver $k$. We consider the Swerling-I sensing model, i.e. the deterministic RCS assumption where RCSs are assumed to be deterministic but unknown during the observation time [48, 38, 40, 41, 49]. The lowpass equivalent of the reflected signal from the $n$th transmitter by a target received at receiver $k$ is given by [^1] The transmitted signals over different transmitters are orthogonal and maintain approximate orthogonality over different delays and Doppler frequencies such that the signals contributed from different transmitters can be separated at each receiver [50, 8, 34, 35, 36].
@@
where $z_{n,k}(t)$ denotes the clutter-plus-noise and $y_{n,k}(t)$ is the loss-free and noise-free received signal,
@@
Note that clutters can initially be mitigated by Doppler processing or space-time adaptive processing [51] and whitened through a whitening filter [34]. It is reasonable to assume the clutter-plus-noise to be additive white Gaussian noise with zero mean and variance $\sigma_{\rm cn}^{2}$ [8, 52, 53, 34, 35, 36]. Define $\delta_{\rm cn}=\frac{P}{T_{\rm eff}\sigma_{\rm cn}^{2}}$ as the signal-to-clutter-plus-noise ratio (SCNR).
The received signal of the communication user is given by,
${r}_{u}(t)$ $=\sqrt{P}\sum_{n=1}^{N}g_{n}{b}_{n}\sqrt{\rho_{n}}s_{n}(t)+{z}(t),$ (6)
where $g_{n}$ is the channel fading between the transmitter $n$ and the user, and the interference-plus-noise ${z}(t)$ is assumed to be the additive circular Gaussian white with zero mean and variance $\sigma_{z}^{2}$. Then, the SINR at the user terminal can be given by
$\gamma$ $=\sum_{n=1}^{N}|g_{n}{b}_{n}|^{2}\rho_{n}\delta\overset{(\text{i})}{=}\delta{\bm{\rho}}^{\rm T}{\bm{g}}$ (7)
where $\delta=\frac{P}{T_{\rm eff}\sigma_{z}^{2}}$ is the total SINR and ${\bm{g}}$ denotes the channel gain [^2] The communication channels are assumed to be estimated a priori using pilot signals [9, 10, 11, 12, 13]. with its $n$th element being $|g_{n}|^{2}$. Note that $(\mathrm{i})$ holds when using a normalized conjugate beamforming weight with ${b}_{n}={g}^{*}_{n}/|g_{n}|$.
III Sensing Metric

### v2

# Discrete-Time CRLB-based Power Allocation for CF MIMO-ISAC with Joint Localization and Velocity Sensing

@@ Abstract
In this paper, we investigate integrated sensing and communication (ISAC) in a cell-free (CF) multiple-input multiple-output (MIMO) network, where each access point functions either as an ISAC transmitter or as a sensing receiver. We devote into the ISAC sensing metric using the discrete-time signal-based Cramér–Rao lower bounds (CRLBs) for joint location and velocity estimation under arbitrary power allocation ratios under the deterministic radar cross section assumption (RCS). Then, we consider the power allocation optimization problem for the CF MIMO-ISAC as the maximization of the communication signal-to-interference-plus-noise ratio (SINR), subject to CRLB-based sensing constraints and per-transmitter power limits. To solve the resulting nonlinear and non-convex problem, we propose a penalty function and projection-based modified conjugate gradient algorithm with inexact line search (PP-MCG-ILS), and an alternative method based on a modified steepest descent approach (PP-MSD-ILS). We show that the proposed algorithms are scalable and can be extended to a broad class of optimization problems involving nonlinear inequality constraints and affine equality constraints. In addition, we extend the PP-MCG-ILS algorithm to the pure sensing scenario, where a penalty function-based normalized conjugate gradient algorithm (P-NCG-ILS) is developed for sensing power minimization. Finally, we analyze the convergence behavior and qualitatively compare the computational complexity of the proposed algorithms. Simulation results confirm the accuracy of the derived CRLBs and demonstrate the effectiveness of the proposed power allocation strategies in enhancing both sensing and overall ISAC performance.
@@ I Introduction
Integrated sensing and communication (ISAC) has emerged as a key enabling technology for sixth-generation (6G) applications, including smart cities, autonomous driving, and remote healthcare [1]. Unlike traditional systems that separate the design of communication and sensing functionalities, ISAC integrates both within a unified framework using shared hardware platforms and spectral resources. This integration offers substantial improvements in spectral efficiency, energy savings, and hardware utilization [2, 3].
ISAC systems can be broadly categorized based on the spatial configuration of transmitters and receivers. Monostatic ISAC systems, where the transmitter and receiver are co-located, typically rely on advanced beamforming techniques to mitigate self-interference in full-duplex operations [4, 5]. In contrast, multistatic ISAC systems deploy spatially separated transmitters and receivers, facilitating spatial diversity and sensing coverage [6, 7]. Building upon the multistatic architecture, the concept of cell-free (CF) multiple-input and multiple-output (MIMO) ISAC has recently emerged [8, 9]. In CF MIMO-ISAC systems, a large number of distributed access points (APs) jointly serve communication users and provide sensing services in a coordinated manner without relying on conventional cell boundaries. This architecture inherits the advantages of both distributed multistatic radar and CF MIMO communication, offering improved spectral efficiency, spatial resolution, and system scalability [10].
Recent advancements in CF MIMO-ISAC have extended from basic trade-off analysis [11] to complex sensing tasks, including target detection [12], and parameter estimation for radar cross-section (RCS) [13], location [14], velocity [15], and angle [16]. For example, [12] studies the fundamental performance tradeoff between the target detection probability and the user’s achievable rate in ISAC systems by deriving the probability of false alarm and the successful probability of detection for the target of interest. [13] provides a unified framework integrating sensing (RCS estimation), computing, and communication to optimize limited system resource. [14] proposes a two-stage scheme for target localization, with communication signals reused as sensing reference signals. Additionally, joint sensing and communication beamforming has been explored in [17].
@@
Despite these advances, a fundamental challenge in ISAC system design, including CF MIMO-ISAC, is the formulation of robust performance metrics, which are critical for guiding system design and balancing communication-sensing trade-offs [2, 3]. While unified metrics based on mutual information [18] and Kullback–Leibler divergence [19] have been proposed for general analysis, adapting these metrics to task-specific sensing objectives remains a significant challenge. Consequently, task-specific metrics for sensing and communication are still widely adopted.
For communication, Shannon capacity underpins metrics such as spectral efficiency [20], achievable rate [21], and signal-to-interference-plus-noise ratio (SINR). Additional metrics, including coverage probability and energy efficiency, have also been considered, particularly in scenarios with stringent quality-of-service (QoS) or energy constraints [22]. On the sensing side, a variety of performance metrics have been investigated, such as sensing mutual information [23], sensing SINR [24, 25], and the Cramér–Rao lower bound (CRLB) [26, 27, 28, 29]. Among these, CRLB-based formulations have proven particularly effective for parameter estimation and algorithm benchmarking. CRLB-based approaches generally fall into two categories depending on the RCS model: deterministic and stochastic. Under the stochastic RCS assumption, RCS values are modeled as random variables, and the corresponding likelihood functions are derived by averaging over their statistical distributions. In contrast, the deterministic RCS assumption treats RCS values as deterministic but unknown quantities that may vary spatially. This model is particularly applicable in slowly varying environments or when prior RCS knowledge is unavailable. CRLBs for both single-target and multiple-target location estimation have been analyzed under the stochastic RCS assumption [26, 27] and the deterministic RCS assumption [28, 29].
@@
However, most existing CRLB formulations are built upon continuous-time signal models, whereas practical radar systems operate in the discrete-time domain, where the sampling rate critically affects sensing accuracy. Our prior work [30] addresses this gap by introducing a discrete-time likelihood function over Gaussian channels, offering a more realistic modeling framework. Building on this foundation, this paper devotes into the CRLBs derivation for joint location and velocity estimation under the deterministic RCS assumption, covering both single- and multi-target cases in CF MIMO-ISAC systems.
@@
Beyond performance characterization, another critical challenge lies in resource allocation, involving time [31], frequency [32], space [33], and power [24, 25, 34, 35, 36]. Recent works in CF MIMO-ISAC systems have addressed power allocation (PA) under sensing SINR constraints [24, 25] or sensing CRLB constraints [34, 35, 36]. These involved sensing CRLBs typically consider localization only, neglecting velocity or angular estimation, and depend on continuous-time models that overlook the impact of sampling rate.
These considerations motivate us to consider a more comprehensive and realistic setting: joint location and velocity estimation in CF MIMO-ISAC systems under a discrete-time signal model with a deterministic RCS assumption. The CRLBs are serve as effective sensing metrics to guide PA across distributed APs, with the goal of maximizing communication SINR. This leads to a class of non-convex and highly nonlinear optimization problems. To address this issue, we develop a tailored optimization framework that integrates penalty methods, projection operations, and gradient-based search, enabling efficient and reliable algorithmic solutions under complex practical constraints. The main contributions are as follows:
  • We derive novel closed-form approximate CRLBs for joint location and velocity estimation under arbitrary PA strategies. These CRLBs are rigorously validated against maximum likelihood estimation (MLE) results, confirming their effectiveness as practical sensing performance metrics. Based on these metrics, we formulate the PA optimization problems for ISAC, aiming at communication SINR maximization while ensuring joint CRLB and power budget constraints.
@@
  • To tackle the non-convexity and nonlinearity in PA optimization for ISAC, we propose a penalty-function and projection-based modified conjugate gradient algorithm with inexact line search (PP-MCG-ILS) together with PP-MSD-ILS algorithm based on the modified steepest descent algorithm (MSD). We further propose a penalty-function-based normalized modified conjugate gradient algorithm with inexact line search (P-NCG-ILS), which can effectively solve the power minimization problem subject to sensing accuracy constraints.
@@
  • We provide a comprehensive analysis of their convergence behavior and computational complexity. The design ensures stable updates, rapid convergence and low computational overhead. Additionally, we show that the proposed PP-MCG-ILS and P-NCG-ILS algorithms can be broadly scalable to a class of constrained non-convex optimization problems involving affine equality constraints (handled via projection) and nonlinear inequality constraints (handled via penalties).
  • Extensive simulations validate the performance of the developed algorithms in both ISAC and sensing-only scenarios. The PP-MCG-ILS algorithm consistently exhibits fast convergence, robust numerical stability, and optimal PA efficiency across various conditions, as compared to benchmark methods. Results also highlight important trade-offs between sensing and communication performance and confirm that the proposed methods can meet stringent sensing QoS while maximizing communication throughput or minimizing power usage.
@@
Section II introduces the received signal model for the CF MIMO-ISAC system. Section III-B formulates the power allocation problem for CF MIMO-ISAC. Section IV presents the proposed power allocation algorithms for both CF MIMO-ISAC and the sensing-only scenario. Section VI analyzes the convergence properties and computational complexity of the proposed algorithms. Section VII provides simulation results to validate the theoretical analysis and algorithmic performance. Finally, Section VIII concludes the paper.
@@ II Signal Model
As shown in Fig. 1, we consider a CF MIMO-ISAC system with widely separated APs. Each AP either serves as an ISAC transmitter or a sensing receiver, similar to the setup in [24]. Assume that there are $N$ single-antenna transmitters and $K$ single-antenna receivers, with the positions of the transmitter $n$ and receiver $k$ denoted as ${\bm{l}}_{n}=[x_{n},y_{n}]^{\rm T}$ and ${\bm{l}}_{k}=[x_{k},y_{k}]^{\rm T}$, respectively. All transmitters and receivers are scheduled and managed by the edge cloud and the core network, e.g., the fifth generation (5G) communication networks. Denote the location and velocity vectors of the sensing target as ${\bm{l}}=[x,y]^{\rm T}$ and ${\bm{v}}=[v_{x},v_{y}]^{\rm T}$, respectively.
@@
The Doppler frequency induced by the relative radial velocity of the target with respect to transmitter $n$ and receiver $k$ is computed as
$f_{n,k}=\dfrac{1}{\lambda}{\bm{v}}^{\rm T}\left(\frac{{\bm{l}}_{n}-{\bm{l}}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}+\frac{{\bm{l}}_{k}-{\bm{l}}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\right),$ (1)
where $\lambda$ denotes the wavelength. Denote $\tau_{n,k}$ as the propagation delay between the $n$th transmitter and the $k$th receiver, as reflected by the target. Then, we have
$\tau_{n,k}$ $=\tau_{n}+\tau_{k},$ (2a)
$\tau_{n}={\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}/{c},$ $\quad\tau_{k}={\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}/{c},$ (2b)
where $\tau_{n}$ and $\tau_{k}$ represent the propagation delays from the $n$th transmitter to the $q$th target, and from the target to the $k$th receiver, respectively, and $c$ denotes the speed of light.
@@
where $P$ denotes the total available energy, $\rho_{n}$ is the PA factor with $\sum_{n=1}^{N}\rho_{n}=1$, ${b}_{n}\in\mathbb{C}$ is the normalized transmit beamforming weight with $|{b}_{n}|^{2}=1$, and $\varsigma\in\mathbb{C}$ has unit variance $\mathbb{E}\{|\varsigma|^{2}\}=1$. Let ${\bm{\rho}}=[\rho_{1},\rho_{2},\cdots,\rho_{N}]^{\rm T}$ be the vector of the PA coefficient of the $N$ transmitters.
Sensing model: For target-reflected reception, we refer the term “RCS” as encompassing both channel fading and radar RCS. Denote the RCS associated with transmitter $n$ and receiver $k$ as ${\alpha}_{n,k}$. We consider the Swerling-I sensing model [37], i.e. the deterministic RCS assumption where RCSs are assumed to be deterministic but unknown during the observation time [29, 28]. Then, the lowpass equivalent of the reflected signal from the $n$th transmitter by a target received at receiver $k$ is given by [^1] The transmitted signals over different transmitters are orthogonal and maintain approximate orthogonality over different delays and Doppler frequencies such that the signals contributed from different transmitters can be separated at each receiver [29, 26].
@@
where $z_{n,k}(t)$ denotes the clutter-plus-noise and $y_{n,k}(t)$ is the loss-free and noise-free received signal, i.e.,
@@
Define the signal-to-clutter-plus-noise ratio (SCNR) as $\delta_{\rm cn}=\frac{P}{T_{\rm eff}\sigma_{\rm cn}^{2}}$ with $T_{\rm eff}$ denoting the effective waveform width.
Communication model: The received signal of the communication user is given by
$\small{r}_{u}(t)$ $=\sqrt{P}\sum\nolimits_{n=1}^{N}g_{n}{b}_{n}\sqrt{\rho_{n}}s_{n}(t)+{z}(t),$ (6)
where $g_{n}$ is the channel fading between the transmitter $n$ and the user, and $z(t)$ denotes the interference-plus-noise with zero mean and variance $\sigma_{z}^{2}$. Then, the SINR at the user terminal can be given by
$\small\gamma$ $=\sum\nolimits_{n=1}^{N}|g_{n}{b}_{n}|^{2}\rho_{n}\delta\overset{(\text{i})}{=}\delta{\bm{\rho}}^{\rm T}{\bm{g}}$ (7)
where $\delta=\frac{P}{T_{\rm eff}\sigma_{z}^{2}}$ is the total SINR and ${\bm{g}}$ is the channel gain [^2] The communication channels are assumed to be estimated a priori using pilot signals [24, 35, 34, 36]. with its $n$th element being $|g_{n}|^{2}$. Note that $(\mathrm{i})$ holds when using a normalized conjugate beamforming with ${b}_{n}={g}^{*}_{n}/|g_{n}|$.
III The Proposed Sensing Metric and Problem Formulation